A window air conditioner consumes 1700W of electric power and has COP = 2.5.

If it runs continuously for 24h , how much heat gets removed from the house?
Answer #1


Applyn the formula for Coeffcient of performance

CoP = Q/W

Cop = Q/t /W/t

Q/t   = = Cop * W/t

Q/t     =   1700 * 2.5

Q /t = 4250 Watts

energy Q = Pt

Q   = 4250 * 24* 60*60

Q = 3.67 *10^8 Joules